OpenStudy (calculusxy):

Jack is twice as old as Lillian. Six years ago Jack was five times as old as Lillian was then. How old is Lillian now? @whpalmer4 @phi @e.mccormick

OpenStudy (anonymous):

Jack is twice as old as Lillian J = 2L Six years ago Jack was five times as old as Lillian was then J-6 = 5(L-6) solve for L

OpenStudy (calculusxy):

Can you help me solve for L? So far i have only got up to: J-6=5L-30

OpenStudy (anonymous):

since J = 2L, replace J with 2L in the equation above

OpenStudy (calculusxy):

2L-6=5L-30 ?

OpenStudy (calculusxy):

@sourwing

OpenStudy (whpalmer4):

Yes, that's right, solve for L to get Lillian's current age.

OpenStudy (calculusxy):

But how can i solve for L? i need help on that.

OpenStudy (whpalmer4):

\[2L - 6 = 5L - 30\]You want to get L alone on one side of the equals sign, and a number on the other. You can do whatever you want, as long as you do it to both sides of the equation. You might start by adding 6 to both sides, for example: \[2L - 6 + 6 = 5L - 30 + 6\]\[2L = 5L -24\]Any ideas?

OpenStudy (calculusxy):

How did u get 24? understood that you subtracted 30-6=24 but why subtract?

OpenStudy (whpalmer4):

No, I didn't subtract, I added -30 + 6 But here, let's do it a different way: \[2L - 6 = 5L - 30\]Let's add 30 to both sides: \[2L - 6 + 30 = 5L - 30 + 30\]\[2L - 6 + 30 = 5L\]Comfortable with that?

OpenStudy (calculusxy):

yes \[2L-24=5L\] now what do i do?

OpenStudy (whpalmer4):

No, -6 + 30 = 30 - 6 =

OpenStudy (calculusxy):

it's okay. now i got the problem.

OpenStudy (calculusxy):

thanks